OpenStudy (anonymous):

Suppose you place bacterium in a jar. Each bacterium divided into 2 bacterium every hour. How much bacterium is in the jar after 7 hours?

OpenStudy (callisto):

 From that , we can know that the no of bacterium is 2^n, where n is the number of hours number of bacteria = 2^n = 2^7 = Can you do it yourself?